5). Calculate the maximum concentration (in M) of silver ions (Ag+) in a solution that contains 0.00125 M of CO32-. The Ksp of Ag2CO3 is 8.46×10-12.
Answer #1

[CO32-] = 0.00125 M

At equilibrium:

Ag2CO3 <----> 2 Ag+ + CO32-

   2s 1.25*10^-3 + s

Ksp = [Ag+]^2[CO32-]

8.46*10^-12=(2s)^2*(1.25*10^-3+ s)

Since Ksp is small, s can be ignored as compared to 1.25*10^-3

Above expression thus becomes:

8.46*10^-12=(2s)^2*(1.25*10^-3)

8.46*10^-12= 4(s)^2 * 1.25*10^-3

s = 4.11*10^-5 M

So,

[Ag+] = 2s

= 2*4.11*10^-5 M

= 8.22*10^-5 M

Answer: 8.22*10^-5 M